# BemaGANv2: A Tutorial and Comparative Survey of GAN-based Vocoders for Long-Term Audio Generation

### v1

@@ Abstract
This paper presents a tutorial-style survey and implementation guide of BemaGANv2, an advanced GAN-based vocoder designed for high-fidelity and long-term audio generation. Built upon the original BemaGAN architecture, BemaGANv2 incorporates major architectural innovations by replacing traditional ResBlocks in the generator with the Anti-aliased Multi-Periodicity composition (AMP) module, which internally applies the Snake activation function to better model periodic structures. In the discriminator framework, we integrate the Multi-Envelope Discriminator (MED), a novel architecture we proposed, to extract rich temporal envelope features crucial for periodicity detection. Coupled with the Multi-Resolution Discriminator (MRD), this combination enables more accurate modeling of long-range dependencies in audio. We systematically evaluate various discriminator configurations, including MSD + MED, MSD + MRD, and MPD + MED + MRD, using objective metrics (FAD, SSIM, PLCC, MCD) and subjective evaluations (MOS, SMOS). This paper also provides a comprehensive tutorial on the model architecture, training methodology, and implementation to promote reproducibility. The code and pre-trained models are available at: https://github.com/dinhoitt/BemaGANv2.
@@ 1 Introduction
In recent years, deep learning-based models for Text-to-Audio (TTA) and Text-to-Music (TTM) Models have attracted considerable interest, in the context of AI-driven music Generation. One of the key components of diffusion-based generative models is the vocoder, which transforms intermediate acoustic representations such as Mel-spectrograms into raw waveform audio signals. The performance and characteristics of the vocoder significantly influence the realism and fidelity of the generated audio, particularly for long-duration outputs.
To address limitations in periodicity modeling and long term dependency, our research group previously introduced BemaGAN[1], a GAN-based vocoder inspired by the HiFi-GAN[2] framework. Notably, BemaGAN replaced the standard Multi-Scale Discriminator (MSD) with a custom-designed Multi-Envelope Discriminator (MED)[1, 3], enabling more sensitive detection of temporal envelope variations to enhance long-term audio quality.
Building upon this foundation, we developed BemaGANv2[3], an improved architecture that incorporates multiple architectural enhancements. These include the integration of the Anti-aliased Multi-Periodicity composition (AMP) module[4] in place of conventional ResBlocks within the generator, and the use of the Snake activation function[5] to strengthen periodic signal modeling. In addition, the frequency coverage was extended from 8kHz to 12kHz with a 24kHz sampling rate, and the discriminator was restructured by introducing the Multi-Resolution Discriminator (MRD) in place of the Multi-Period Discriminator (MPD).
This paper serves two primary purposes:
@@
  • Provide a tutorial-style explanation of the architectural principles, training methodology, and implementation details of BemaGANv2.
  • Validate the effectiveness of the discriminator combination used in BemaGANv2 by conducting a comparative analysis of various discriminator configurations, including MSD+MED, MSD+MRD, and MPD+MED+MRD, and evaluate their performance using both objective and subjective metrics.
  • By consolidating the architectural insights and experimental outcomes, this paper aims to facilitate understanding and reproducibility of advanced vocoder systems, and to support further research on long-term, high-fidelity audio generation.
@@
All experiments were conducted using the LJSpeech dataset for training and random audio samples from freesound.org for testing, on an NVIDIA A100 GPU environment via Google Colab.
This article is an extended version of our ICAIIC 2025 paper[3], enriched with additional experiments, theoretical and mathematical analysis, and presented as a survey and tutorial to enhance clarity and reproducibility.
@@ 2.1 Overview of Vocoders
In TTA and TTM systems, the vocoder plays a critical role in converting intermediate acoustic representations—typically Mel-spectrograms—into time-domain audio waveforms. The quality and structure of the vocoder significantly affect the fidelity, naturalness, and temporal coherence of the generated audio. In fact, most diffusion-based TTA and TTM models, such as AudioLDM[6, 7],FLUX Music[8], commonly adopt HiFi-GAN as the vocoder due to its high-quality reconstruction and efficient inference. In this subsection, we review key developments in parallel GAN-based vocoders, focusing on MelGAN[9], HiFi-GAN, and BigVGAN, to illustrate the evolution of architecture design and its implications for periodicity modeling.
@@ 2.1.2 HiFi-GAN
Model Generator Discriminator(s) Activation Key Features
MelGAN Transposed Conv + ResBlock MSD Leaky ReLU Fast inference, low fidelity
HiFi-GAN MRF-based ResBlocks MPD + MSD Leaky ReLU Harmonic modeling, MRF design
BigVGAN AMP block MPD + MRD Snake Anti-aliasing, large-scale training
BemaGAN HiFi-GAN-style ResBlock MED + MPD Leaky ReLU Envelope-aware discrimination
BemaGANv2 AMP block MED + MRD Snake Balanced fidelity and periodicity
@@ 2.1.3 BigVGAN
BigVGAN is a universal GAN-based vocoder designed to synthesize high-fidelity audio across a wide range of conditions, including unseen speakers, languages, and noisy recording environments. It builds upon the generator architecture of HiFi-GAN but introduces several architectural and training innovations to address the limitations of prior models in capturing periodicity, extrapolation, and robustness to out-of-distribution (OOD) inputs[5].
The core architectural contribution of BigVGAN is the AMP (Anti-aliased Multi-Periodicity) block, which replaces the traditional residual blocks in the generator. Each AMP block consists of a series of dilated convolutions embedded with channel-wise Snake activation functions, thereby enforcing an explicit periodic inductive bias in the transformation process. Each AMP block contains a series of dilated convolutions with channel-wise Snake activation functions, which provide an explicit periodic inductive bias. The Snake function, defined as:
@@
injects learnable periodicity into the activation, enabling the generator to better model complex harmonic structures in audio. This significantly improves the model’s ability to generalize and extrapolate periodic content, which is particularly beneficial in musical or expressive vocal signals[4].
@@ 3 Model
In this section, we provide a detailed tutorial-style description of the architectural components of BemaGANv2. The model adopts a GAN-based vocoder framework, integrating periodicity-aware generator and perceptually motivated discriminator. We begin by describing the generator based on AMP blocks, followed by the Multi-Envelope Discriminator (MED) and Multi-Resolution Discriminator (MRD), and conclude with the training objectives.
@@ 3.1 Generator with AMP Block
The generator in BemaGANv2 is structured similarly to HiFi-GAN, using transposed convolution layers for upsampling and residual blocks for waveform synthesis. However, instead of conventional residual blocks, it employs the AMP block, originally introduced in BigVGAN[4].
Each AMP block contains dilated convolutions paired with the Snake activation function, which provides a learnable periodic bias. This allows the generator to better model harmonic structures and rhythmic patterns in long-term audio generation. The overall design enhances both local detail and global periodic consistency[4][5].
@@ 3.2 Multi-Envelope Discriminator (MED)
A central contribution of BemaGANv2 is the introduction of the Multi-Envelope Discriminator (MED)[1][3], a novel discriminator architecture proposed by our research group. Rather than operating directly on raw waveforms or spectrograms, Multi-Envelope Discriminator (MED) analyzes time-domain envelopes derived from the audio signal. These envelopes are extracted using multiple low-pass filters with varying cutoff frequencies, capturing energy dynamics at different temporal scales.
Each envelope is processed by a dedicated 1D convolutional stack, and the outputs are aggregated to form the discriminator output. This design enables the model to detect patterns related to phrasing, prosody, and amplitude modulation properties closely linked to human perception of naturalness. Multi-Envelope Discriminator (MED) thus plays a crucial role in enhancing long-term temporal coherence in the generated audio.
@@ 3.3 Multi-Resolution Discriminator (MRD)
To complement Multi-Envelope Discriminator (MED), BemaGANv2 also incorporates a Multi-Resolution Discriminator (MRD). This module operates in the time-frequency domain using STFT-based log-magnitude spectrograms. By processing multiple spectrograms computed with different FFT sizes and hop lengths, Multi-Resolution Discriminator (MRD) enforces spectral consistency across various resolutions.
The Multi-Resolution Discriminator (MRD) primarily helps the model maintain harmonic sharpness and timbral accuracy, which are important for high-fidelity generation. Together with Multi-Envelope Discriminator (MED), Multi-Resolution Discriminator (MRD) provides complementary perspectives temporal envelope and spectral structure guiding the generator to produce perceptually convincing waveforms.
@@ 3.4 Overall Architecture
Figure 3 presents the complete BemaGANv2 architecture. The generator transforms the input Mel-spectrogram into a waveform using AMP blocks. The output waveform is then assessed by both Multi-Envelope Discriminator (MED) and Multi-Resolution Discriminator (MRD), which provide feedback based on temporal energy flow and spectral consistency, respectively. This dual-discriminator setup enables BemaGANv2 to generate audio that is both rhythmically natural and spectrally accurate.
@@ 3.5 Training Loss
where $K$ is the number of sub-discriminators, and $\lambda_{fm},\lambda_{mel}$ are hyperparameters for feature matching and mel-spectrogram losses, respectively.

### v3

# BemaGANv2: Discriminator Combination Strategies for GAN-based Vocoders in Long-Term Audio Generation

@@ Abstract
This paper presents BemaGANv2, an advanced GAN-based vocoder designed for high-fidelity and long-term audio generation, with a focus on systematic evaluation of discriminator combination strategies. Long-term audio generation is critical for applications in Text-to-Music (TTM) and Text-to-Audio (TTA) systems, where maintaining temporal coherence, prosodic consistency, and harmonic structure over extended durations remains a significant challenge. Built upon the original BemaGAN architecture, BemaGANv2 incorporates major architectural innovations by replacing traditional ResBlocks in the generator with the Anti-aliased Multi-Periodicity composition (AMP) module, which internally applies the Snake activation function to better model periodic structures. In the discriminator framework, we integrate the Multi-Envelope Discriminator (MED), a novel architecture we proposed, to extract rich temporal envelope features crucial for periodicity detection. Coupled with the Multi-Resolution Discriminator (MRD), this combination enables more accurate modeling of long-range dependencies in audio. We systematically evaluate various discriminator configurations, including Multi-Scale Discriminator (MSD) + MED, MSD + MRD, and Multi-Period Discriminator (MPD) + MED + MRD, using objective metrics (Fréchet Audio Distance (FAD), Structural Similarity Index (SSIM), Pearson Correlation Coefficient (PCC), Mel-Cepstral Distortion (MCD), Multi-Resolution STFT (M-STFT), Periodicity error (Periodicity)) and subjective evaluations (MOS, SMOS). To support reproducibility, we provide detailed architectural descriptions, training configurations, and complete implementation details. The code, pre-trained models, and audio demo samples are available at: https://github.com/dinhoitt/BemaGANv2.
@@ 1 Introduction
In recent years, deep learning-based generative models for Text-to-Audio (TTA) and Text-to-Music (TTM) have gained significant attention due to their capability to produce high-fidelity audio. In diffusion-based architectures, the vocoder serves as a critical component that transforms intermediate acoustic representations such as Mel-spectrograms into time-domain waveform signals. The quality and structure of the vocoder directly influence the realism and continuity of the generated audio, particularly in long-form generation tasks.
Early models such as MelGAN [1] emphasized inference efficiency through fully convolutional architectures but struggled with harmonic fidelity. HiFi-GAN [2] addressed this by introducing a Multi-Period Discriminator (MPD) that improved periodicity modeling. Subsequently, BigVGAN extended this line of work with the Anti-aliased Multi-Periodicity (AMP) module and the Snake activation [3, 4] to better capture harmonic structures and long-term coherence.
Our previous work, BemaGAN [5], introduced the Multi-Envelope Discriminator (MED), which focuses on capturing temporal energy envelopes through both Hilbert transform-based envelope extraction (for upper and lower bounds) and low-pass filtered signals (for multi-scale amplitude variations ). Building on this, BemaGANv2 combines several architectural enhancements: the AMP block replaces the ResBlock in the generator; the Snake activation is employed to improve periodic signal modeling; and the discriminator is extended with the Multi-Resolution Discriminator (MRD) to refine spectral consistency [6]. Detailed descriptions of the AMP block, Snake activation, and MRD are provided in Section 3. While diffusion-based vocoders such as WaveGrad [7] and DiffWave [8] have demonstrated impressive audio quality, their iterative sampling process increases computational cost and latency, making them less suitable for long-term generation and real-time streaming applications. Therefore, this work focuses exclusively on GAN-based vocoders, which offer fast, single-pass inference while maintaining high fidelity—a critical requirement for deployment in TTA and TTM systems.
This paper provides a tutorial-style study of GAN-based vocoders with a focus on discriminator design for long-term audio generation. While diffusion-based, autoregressive, and flow-based vocoders represent important parallel research directions, we limit our scope to GAN-based models to maintain coherence and alignment with long-term generation and real-time streaming requirements. Specifically, our goals are threefold:
@@
  • To provide a tutorial-style overview of the core architectural principles, training methodology, and implementation strategies of modern GAN-based vocoders.
  • To systematically analyze how different discriminator combinations affect audio reconstruction quality under identical generator and training conditions, demonstrating that the choice of discriminator combination is a critical design factor.
  • To present BemaGANv2 (MED+MRD) as a case study demonstrating that complementary discriminator pairing—temporal envelope modeling via MED and spectral consistency via MRD—achieves the most balanced performance, rather than simply replacing any single existing discriminator.
@@
All experiments discussed herein were conducted using the LJSpeech dataset for training and freeform audio from Freesound.org for evaluation. Computation was performed on an NVIDIA A100 GPU via Google Colab.
LJSpeech Dataset [20]: A widely-used single-speaker English speech corpus consisting of 13,100 short audio clips with a total duration of approximately 24 hours. Recorded by professional speaker Linda Johnson, the dataset features clean, high-quality 16-bit PCM audio at 22,050 Hz sampling rate. LJSpeech has become a standard benchmark for speech synthesis and vocoder evaluation due to its consistent recording quality and phonetic diversity.
Freesound.org Audio: A collaborative database of Creative Commons-licensed audio samples spanning diverse acoustic conditions, including environmental sounds, musical instruments, sound effects, and ambient recordings. We selected samples from Freesound.org to evaluate model generalization beyond speech, testing vocoder performance on out-of-distribution audio with complex spectral characteristics and long-term temporal structures.
This paper is an extended version of our prior work published at ICAIIC 2025 [6], enriched with additional experiments, comparative discussions, and theoretical analysis. The content has been restructured to support reproducibility, enriched with ablation experiments, expanded evaluation metrics, and theoretical analysis.
@@ 2.1 Overview of Vocoders
In TTA and TTM systems, the vocoder plays a critical role in converting intermediate acoustic representations—typically Mel-spectrograms—into time-domain audio waveforms. The quality and structure of the vocoder significantly affect the fidelity, naturalness, and temporal coherence of the generated audio. In fact, most diffusion-based TTA and TTM models, such as AudioLDM[9, 10],FLUX Music[11], commonly adopt HiFi-GAN as the vocoder due to its high-quality reconstruction and efficient inference. In this subsection, we review key developments in parallel GAN-based vocoders, focusing on MelGAN, HiFi-GAN, and BigVGAN, to illustrate the evolution of architecture design and its implications for periodicity modeling. Following this, we also discuss our prior work, BemaGAN , and its advanced iteration, BemaGANv2, which serve as a culmination of these architectural advancements.
@@ 2.1.2 HiFi-GAN
Model Generator Discriminators Activation Key Features Limitations
MelGAN Transposed Conv + ResBlock MSD Leaky ReLU Real-time inference Poor periodicity
HiFi-GAN MRF-based ResBlocks MPD + MSD Leaky ReLU Harmonic modeling Fixed periodic hops
BigVGAN AMP blocks MPD + MRD Snake Anti-aliasing, expressive periodicity High complexity
BemaGAN HiFi-GAN ResBlock MPD + MED Leaky ReLU Envelope-aware detection Generator underpowered
BemaGANv2 AMP block MED + MRD Snake Balanced temporal and spectral modeling Single-speaker training data
@@ 2.1.3 BigVGAN
BigVGAN is a universal GAN-based vocoder designed to synthesize high-fidelity audio across a wide range of conditions, including unseen speakers, languages, and noisy recording environments. It builds upon the generator architecture of HiFi-GAN but introduces several architectural and training innovations to address the limitations of prior models in capturing periodicity, extrapolation, and robustness to out-of-distribution (OOD) inputs[3, 4].
The core architectural contribution of BigVGAN is the AMP (Anti-aliased Multi-Periodicity) block, which replaces the traditional residual blocks in the generator. Each AMP block embeds channel-wise Snake activation functions within a series of dilated convolutions, thereby enforcing an explicit periodic inductive bias in the transformation process. Each AMP block contains a series of dilated convolutions with channel-wise Snake activation functions, which provide an explicit periodic inductive bias. The Snake function, defined as:
@@
where α is a learnable frequency parameter that controls the periodicity of the activation. A higher α increases the frequency of oscillation, allowing the network to model fine-grained harmonic details, while a lower α captures broader periodic structures. In our implementation, α is initialized randomly and optimized during training via backpropagation.
This learnable periodicity significantly improves the model’s ability to generalize and extrapolate periodic content, which is particularly beneficial in musical or expressive vocal signals[3]
@@ 2.2 Snake Function and Periodictiy in Audio
5-505502244xxf(x)f(x)ReLU5-505502244xxf(x)f(x)Leaky ReLU5-50551-1011xxf(x)f(x)Tanh5-50551-1011xxf(x)f(x)Sin5-50555-5055xxf(x)f(x)Snake ($\alpha=1$)
Figure 1: Comparison of activation functions (each in separate axis).
@@ 3 Model
In this section, we provide a detailed tutorial-style description of the architectural components of BemaGANv2. The model adopts a GAN-based vocoder framework, integrating periodicity-aware generator and perceptually motivated discriminator. We begin by describing the generator based on AMP blocks, followed by the Multi-Envelope Discriminator (MED) and Multi-Resolution Discriminator (MRD), and conclude with the training objectives. This structured approach aims to provide a clear understanding of how each component contributes to BemaGANv2’s overall performance in high-fidelity, long-term audio generation.
@@ 3.1 Generator with AMP Block
The generator in BemaGANv2 is structured similarly to HiFi-GAN, utilizing transposed convolution layers for upsampling and residual blocks for waveform synthesis. However, a key distinction is its employment of the Anti-aliased Multi-Periodicity (AMP) block, originally introduced in BigVGAN[3], which replaces conventional residual blocks.
Each AMP block, depicted in Figure 2’s detailed view, consists of an upsampling layer with Low-Pass Filtering (LPF), followed by a Snake activation function, a downsampling layer with LPF, and dilated convolutions. The LPFs are crucial for anti-aliasing, preventing high-frequency artifacts that can arise from non-linear operations like upsampling and the Snake activation. Each AMP block contains dilated convolutions paired with the Snake activation function, which provides a learnable periodic bias. This allows the generator to better model harmonic structures and rhythmic patterns in long-term audio generation by injecting learnable periodicity into the activation, which is particularly beneficial in musical or expressive vocal signals. The overall design enhances both local detail and global periodic consistency.[3, 4].
@@ 3.2 Multi-Envelope Discriminator (MED)
A central contribution of BemaGANv2 is the introduction of the Multi-Envelope Discriminator (MED)[5, 6], a novel discriminator architecture proposed by our research group. Unlike traditional discriminators that operate directly on raw waveforms or spectrograms, the MED analyzes time-domain envelopes derived from the audio signal. This distinct approach allows it to focus on temporal energy patterns crucial for perceptual quality, such as prosodic variation.
Algorithm 1 Multi-Envelope Discriminator (MED) Forward Pass
1:Input: Real audio $y$, Generated audio $\hat{y}$
2:Output: Discriminator outputs and feature maps
3:Initialize: Envelope extraction modes $M=\{-1,0,1,300,500\}$
4:     1-1: Lower envelope, 0: Original signal, 11: Upper envelope
5:     $300,500$: Butterworth low-pass filtered envelopes (Hz)
[…]
8:for each $m\in M$ do
[…]
12: Append $out_{r}$ to $D_{outputs}^{real}$
13: Append $out_{g}$ to $D_{outputs}^{gen}$
14: Append $fmap_{r}$ to $F_{maps}^{real}$
15: Append $fmap_{g}$ to $F_{maps}^{gen}$
16:end for
[…]
These envelopes are extracted using the Envelope() function with configurations $F=\{-1,0,1,300,500\}$. For $f_{\max}\in\{-1,1\}$, Hilbert transform-based lower and upper envelopes are computed as $|\mathcal{H}(-x(t))|$ and $|\mathcal{H}(x(t))|$, respectively, providing strict amplitude bounds. For $f_{\max}=0$, the original signal is preserved as a baseline. For $f_{\max}\in\{300,500\}$ Hz, Butterworth low-pass filtering precedes envelope extraction, capturing smoothed amplitude variations at syllabic (300 Hz) and phonemic (500 Hz) temporal scales. This multi-scale approach enables the discriminator to evaluate both mathematically rigorous envelope accuracy and perceptually relevant temporal patterns. The use of multiple cutoff frequencies enables a multi-scale analysis of the audio’s temporal structure, providing richer information about phrasing and amplitude modulation. As shown in Figure 3, each extracted envelope is then processed by a dedicated 1D convolutional stack. The outputs from these stacks are subsequently aggregated to form the discriminator’s final output. This design enables the model to detect patterns related to phrasing, prosody, and amplitude modulation properties, which are closely linked to human perception of naturalness. Consequently, the MED plays a crucial role in enhancing long-term temporal coherence in the generated audio
Algorithm 2 Envelope Extraction Methods
1:Input: Audio signal $x$, Envelope mode $m$
2:Output: Extracted envelope signal
3:if m=1m=-1 then
[…]
5:else if $m=0$ then
[…]
7:else if $m=1$ then
[…]
9:else if $m=300$ or $m=500$ then
[…]
12:end if
13:envelopeenvelope
The five envelope extraction modes serve distinct purposes in capturing temporal dynamics:
  • Lower Envelope (m=1m=-1): Computed as the negative of the absolute Hilbert transform, tracking minimum amplitude variations and capturing the lower bound of signal energy.
  • Original Signal ($m=0$): The unfiltered input signal serves as a baseline, providing reference discrimination without envelope-based preprocessing.
  • Upper Envelope ($m=1$): The instantaneous amplitude obtained via absolute Hilbert transform, tracking maximum amplitude variations and representing the upper bound of signal energy.
  • 300 Hz Filtered Envelope ($m=300$): A Butterworth low-pass filter at 300 Hz isolates low-frequency components, with the envelope extracted via Hilbert transform to capture slow prosodic variations.
  • 500 Hz Filtered Envelope ($m=500$): A Butterworth low-pass filter at 500 Hz captures mid-low frequency dynamics, enabling discrimination of intermediate temporal patterns between phoneme-level and syllable-level structures.
This multi-scale envelope analysis enables the MED to detect subtle temporal characteristics across different frequency ranges, enhancing its sensitivity to prosodic variation, phrasing, and amplitude modulation—all of which are crucial for perceptual naturalness in long-term audio generation.
@@ 3.3 Multi-Resolution Discriminator (MRD)
To complement Multi-Envelope Discriminator (MED), BemaGANv2 also incorporates a Multi-Resolution Discriminator (MRD). This module operates in the time-frequency domain using STFT-based log-magnitude spectrograms. By processing multiple spectrograms computed with different FFT sizes and hop lengths, Multi-Resolution Discriminator (MRD) enforces spectral consistency across various resolutions. This multi-resolution analysis allows the discriminator to effectively capture both low-frequency pitch accuracy and high-frequency timbral details.
The Multi-Resolution Discriminator (MRD) primarily helps the model maintain harmonic sharpness and timbral accuracy, which are important for high-fidelity generation. Together with the MED, the MRD provides complementary perspectives—temporal envelope and spectral structure, respectively—thereby guiding the generator to produce perceptually convincing waveforms that are both rhythmically natural and spectrally accurate.
@@ 3.4 Overall Architecture
Figure 4 presents the complete BemaGANv2 architecture. The generator transforms the input Mel-spectrogram into a raw waveform using AMP blocks. The output waveform is then assessed by both Multi-Envelope Discriminator (MED) and Multi-Resolution Discriminator (MRD). The Multi-Envelope Discriminator (MED) provides feedback based on temporal energy flow, focusing on characteristics like phrasing and prosody, while the Multi-Resolution Discriminator (MRD) provides feedback based on spectral consistency across various resolutions, ensuring harmonic sharpness and timbral accuracy. This dual-discriminator setup enables BemaGANv2 to generate audio that is both rhythmically natural and spectrally accurate. Adversarial and auxiliary losses are backpropagated from both discriminators to train the generator.
@@ 3.5 Training Loss
where:
  • $x$ denotes real audio waveforms sampled from the training distribution;
  • $s$ denotes input conditioning (Mel-spectrograms);
  • $G(s)$ represents generated audio waveforms;
  • $D(\cdot)$ represents the discriminator output;
  • $\mathbb{E}_{x,s}[\cdot]$ denotes expectation over the joint distribution of real audio and conditioning.
@@
where $K$ is the number of sub-discriminators, and $\lambda_{fm},\lambda_{mel}$ are hyperparameters for feature matching and mel-spectrogram losses, respectively. These weights balance the contributions of the adversarial, feature matching, and mel-spectrogram losses, guiding the generator to synthesize high-quality and perceptually accurate audio.
In our implementation, we follow the HiFi-GAN protocol with the following loss weights:
  • $\lambda_{\mathrm{fm}}=2$  (feature matching loss weight)
  • $\lambda_{\mathrm{mel}}=45$  (mel-spectrogram reconstruction loss weight)
These values are identical to those used in HiFi-GAN [2] and have been validated across multiple GAN-based vocoder architectures. Complete training hyperparameters are provided in D.2.
Our loss formulation is identical to HiFi-GAN[2], which adopts the Least-Squares GAN (LSGAN) objective[14] for improved training stability. The key difference lies in the discriminator architecture (MED+MRD in BemaGANv2 vs. MPD+MSD in HiFi-GAN), while the loss weighting scheme remains unchanged